OpenStudy (anonymous):

4,3 greatest common factor and least common factor please

OpenStudy (anonymous):

aha! 4 and 3 LCM is 12

OpenStudy (anonymous):

4=2*2*1 3=3*1 GCF=1 LCM=2*2*3=12

OpenStudy (anonymous):

With consecutive numbers, just multiply them together to get LCM

OpenStudy (thepwner):

12